OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (2, 2) and (4, 8)?

OpenStudy (anonymous):

I'm doing a review, and I'd appreciate some help!

OpenStudy (goformit100):

sure

OpenStudy (anonymous):

(y-2)=(8-2)(x-2) =>y-2=6x-12 =>6x-y-10=0

OpenStudy (anonymous):

sorry a typo!!

OpenStudy (anonymous):

(y-2)=(6/2)(x-2) =>y-6=3x-6 =>y=3x

OpenStudy (anonymous):

I'm sorry, I don't understand. Is this point-slope form?